what has high abundance : ArCl or NaCl or KCl

arrange elements to their ionization energy least to greatest : Ar , Xe , Ne , Rn

what is maximum value of ms according to pauli in oxygen atom 1s2 2s2 2p4 : 1 or 2 or 4
Answer #1

1. NaCl has the highest abundance.

2. According to periodic table, the ionization energy Ar= 1520, Xe= 1170, Ne= 2081 and Rn= 1040.

So, least to greatest ionization energy will be Rn<Xe<Ar<Ne

3. The maximum value of ms according to Pauli's exclusion principle will be = 4

Because

for 1s orbital,  n=1, l=0, ml=0, ms=-1/2

and n=1, l=0, ml=0, ms=+1/2

similar for 2s orbital

n=2, l=0, ml=0, ms=-1/2

and n=2, l=0, ml=0, ms=+1/2

and for 2 px

n=2, l=1, ml=-1, ms=-1/2

n=2, l=1, ml=-1, ms=+1/2

and for 2 py

n=2, l=1, ml=0, ms=-1/2

n=2, l=1, ml=0, ms=+1/2